OpenStudy (anonymous):

Single Pules , sine integral find the fourier integral representation of the function

OpenStudy (anonymous):



OpenStudy (anonymous):

i suck at this x_x

OpenStudy (anonymous):

ermm

OpenStudy (anonymous):

i like ur side dark in that pic c:

OpenStudy (anonymous):

^^

OpenStudy (anonymous):

i think @SithsAndGiggles might know something about these or @dan815

OpenStudy (anonymous):

\[F(f) = \int\limits_{-\infty}^{+\infty}f(t) \sin(2 \pi f t)dt = \int\limits_{-1}^{1}1.\sin(2 \pi f t) dt\]

OpenStudy (anonymous):

aha ohk then ?

OpenStudy (anonymous):

calculate that nursery integral -_-

OpenStudy (anonymous):

xD okii

hartnn (hartnn):

sin is an odd function, so from -1 to +1, you can directly write it as 0

hartnn (hartnn):

thats fourier sine integral, which is one of the parts of fourier integral

hartnn (hartnn):

calculate fourier cosine integral too

OpenStudy (anonymous):

ok ill take a snap for my sol

OpenStudy (anonymous):

but the question asks for only sine integral ? :-/

OpenStudy (anonymous):

mm no mashy its directly asking for find the fourier integral representation of the function :D Single Pules , sine integral is the problem name xD

hartnn (hartnn):

if only sine integral is asked, then its representation would be 0 function

OpenStudy (anonymous):

no its ask for fourier

OpenStudy (anonymous):

i think i have an idea how to solve it nw , th guys ill post when i done :D